# Self-Exploring Language Models: Active Preference Elicitation for Online Alignment

### v1

@@ 5.1 Experiment Setup
We select the UltraFeedback [11] dataset as our training set, which contains 61k preference pairs of single-turn conversations. For the ranker providing AI feedback during online alignment, we choose the small-sized PairRM (0.4B) [25]. All experiments are conducted on 8xA100 GPUs.
Due to the absence of performant open-source online direct alignment codebases at the time of this study, we first implement an iterative version of DPO as the baseline, adhering to the same steps as Algorithm 1 but training the LLM with the standard DPO objective. Then we conduct a grid search over hyperparameters, such as the batch size, learning rate, and iteration number, to identify the optimal settings for the iterative DPO baseline. Details on the hyperparameters and grid search results are provided in Appendix C. We follow these best settings to train SELM for a fair comparison. In addition, the top choice for the base models of SELM are LLMs that are finetuned with RLHF after SFT, since they are typically more capable than SFT-only and pertrained models. We consider two series of LLMs: Zephyr [57] and Llama-3 [38], to demonstrate the robustness of SELM. Since the official Zephyr-7B-$\beta$ model is finetuned with DPO on the same UltraFeedback dataset, to avoid overoptimization, we choose Zephyr-7B-SFT [^1] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta as the base model and perform $3$ iterations of SELM after a single iteration of standard DPO training on the first portion of the training data (we refer to this model as Zephyr-7B-DPO). For Llama-3-8B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct that is already finetuned with RLHF, we directly apply $3$ iterations of SELM training.
@@ 5.2 Experiment Results
We first report the performance of SELM and the baselines on the instruction-following chat benchmarks AlpacaEval 2.0 [14] and MT-Bench [71] in Table 1. We can observe that for AlpacaEval 2.0, SELM significantly boosts Zephyr-7B-SFT and Llama-3-8B-Instruct, achieving length-controlled (LC) win rate improvements of $+16.24\%$ and $+11.75\%$, respectively. This enhancement results in models that are competitive with or even superior to much larger LLMs, such as Yi-34B-Chat [65] and Llama-3-70B-Instruct. For the multi-turn MT-Bench, which exhibits higher variance, we report the average scores of SELM and DPO baselines across $3$ runs. We observe that SELM improves the scores by $+2.31$ and $+0.32$, respectively. Furthermore, the proposed method self-explores and enhances the model monotonically, with consistent performance improvements in each iteration. This validates the robustness of our algorithm. Compared to other iterative post-training algorithms, such as SPIN [7], DNO [50], and SPPO [61], SELM gains more improvements on both benchmarks when using the weaker base model (Zephyr-7B-SFT), and achieves the best performance when using Llama-3-8B-Instruct as the base model.
Figure 2: Pairwise length-controlled win rates comparison between SELM, iterative DPO, and base models on the AlpacaEval 2.0 benchmark. Scores represent the LC win rates of the row models against the column models. Models positioned in higher rows have higher LC win rates against the base model and thus better performance.
@@
Table 1: Results on AlpacaEval 2.0 and MT-Bench. Names inside the brackets are the base models that are aligned based upon. The red arrows indicate the increment or decrement from the base model. Compared to iterative DPO and other online alignment baselines, SELM gains more improvements based on the weaker Zephyr-7B-SFT model and achieves superior performance that is competitive with much larger SOTA models when finetuned from Llama-3-8B-Instruct.
@@
Beyond instruction-following benchmarks, we also evaluate SELM and the baselines on several academic benchmarks, including GSM8K [10], HellaSwag [68], ARC challenge [9], TruthfulQA [33], EQ-Bench [45], and OpenBookQA (OBQA) [39]. To better reflect the capabilities of LLMs, we adopt various settings for these benchmarks, including zero-shot, few-shot, and few-shot Chain-of-Thought (CoT) settings. The accuracy results for these multiple-choice QA benchmarks are provided in Table 2. It can be observed that both our method and the baselines can degrade after the RLHF phase on some benchmarks, which is known as the alignment tax [2; 40; 30]. Nevertheless, our method is still able to improve the base models on most of the benchmarks and offers the best overall performance.
We note that SELM is one of the instantiations of the proposed self-exploration objective in (1.1), with reparameterized reward functions and algorithm-specific designs described in Section 3.2, such as the dataset partition and update rule. However, this objective is not restricted to the current implementation and can also be directly applied to any other online alignment framework, with or without a separate reward model, regardless of differences in algorithm designs. Thus, the proposed method is orthogonal to and can be integrated directly into the recent online RLHF workflows [13; 62; 23] that incorporate additional delicate designs with carefully curated datasets.
@@ 5.3 Ablation Study
Next, we study the difference in reward distributions with varying $\alpha$ and iterations. Specifically, we greedily sample from the LLM using prompts from the holdout test set (2k in total) of UltraFeedback and generate rewards for these responses with PairRM. We then calculate the fraction of data that lies in each partition of reward values. The results for different $\alpha$ values of SELM Iter 2 (Zephyr) are shown in Figure 3 (Middle), which indicate that increasing $\alpha$ results in distributions that are more concentrated in higher-reward regions.
@@ 6 Conclusion & Future Work
In this paper, we introduced an active preference elicitation method for the online alignment of large language models. By incorporating an optimism term into the reward-fitting objective, the proposed bilevel self-exploring objective effectively balances between exploiting observed data and exploring potentially high-reward regions. Unlike standard online RLHF algorithms that passively explore the response space by sampling from the training LLM, whose sole objective is maximizing the expected learned reward, our method actively seeks diverse and high-quality responses. This self-exploration mechanism helps mitigate the risk of premature convergence and overfitting when the reward model is only locally accurate. To optimize this bilevel objective, we solve the inner-level problem and reparameterize the reward with the LLM policy, resulting in a simple yet novel iterative alignment algorithm called Self-Exploring Language Models (SELM). Compared to DPO, SELM improves the exploration efficiency by selectively favoring responses with high potential rewards rather than indiscriminately sampling unseen responses.
Our experiments, conducted with Zephyr-7B-SFT and Llama-3-8B-Instruct models, demonstrated the efficacy of SELM. Finetuning on the UltraFeedback dataset and leveraging PairRM for AI feedback, SELM achieved substantial improvements in performance on AlpacaEval 2.0, MT-Bench, and academic benchmarks. These results underscore the ability of SELM to enhance the alignment and capabilities of large language models by promoting more diverse and high-quality responses. Since the proposed technique is orthogonal to the adopted online RLHF workflow, it will be interesting to apply our method within more sophisticated alignment frameworks with advanced designs, which we would like to leave as future work.
@@
For the derivation of the DPO gradient $\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\mathcal{D}_{t})$, we refer the readers to [49].
@@
Appendix C Experiment Setup

### v3

@@ 6.1 Experiment Setup
We adopt UltraFeedback (Cui et al., 2023) as our training dataset, which contains 61k preference pairs of single-turn conversations. For the external ranker during online alignment, we choose the small-sized PairRM (0.4B) (Jiang et al., 2023). All experiments are conducted on 8xA100 GPUs.
Due to the absence of performant open-source online direct alignment codebases at the time of this study, we first implement an iterative version of DPO as the baseline, adhering to the same steps as Algorithm 1 but training the LLM with the standard DPO objective. Then we conduct a grid search over hyperparameters, such as the batch size, learning rate, and iteration number, to identify the optimal settings for the iterative DPO baseline. We follow these best settings to train SELM. In addition, we apply iterative DPO and SELM on instruction fine-tuned models. Specifically, we consider two series of LLMs: Zephyr (Tunstall et al., 2023b) and Llama-3 (Meta, 2024), to demonstrate the robustness of SELM. Since the official Zephyr-7B-$\beta$ model is fine-tuned with DPO on the same UltraFeedback dataset, to avoid overoptimization, we choose Zephyr-7B-SFT [^1] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta as the base model and perform $3$ iterations of SELM after a single iteration of standard DPO training on the first portion of the training data (we refer to this model as Zephyr-7B-DPO). For Llama-3-8B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct that is already fine-tuned with RLHF, we directly apply $3$ iterations of SELM training.
@@ 6.2 Experiment Results
We first report the performance of SELM and the baselines on the instruction-following chat benchmarks AlpacaEval 2.0 (Dubois et al., 2024) and MT-Bench (Zheng et al., 2024) in Table 1. We can observe that for AlpacaEval 2.0, SELM significantly boosts Zephyr-7B-SFT and Llama-3-8B-Instruct, achieving length-controlled (LC) win rate improvements of $+16.24\%$ and $+11.75\%$, respectively. This enhancement results in models that are competitive with or even superior to much larger LLMs, such as Yi-34B-Chat (Young et al., 2024) and Llama-3-70B-Instruct. For the multi-turn MT-Bench, which exhibits higher variance, we report the average scores of SELM and DPO baselines across $3$ runs. We observe that SELM improves the scores by $+2.31$ and $+0.32$, respectively. Furthermore, the proposed method self-explores and enhances the model monotonically, with consistent performance improvements in each iteration. This validates the robustness of our algorithm. Compared to other iterative post-training algorithms, such as SPIN (Chen et al., 2024), DNO (Rosset et al., 2024), and SPPO (Wu et al., 2024), SELM gains more improvements on both benchmarks when using the weaker base model (Zephyr-7B-SFT), and achieves the best performance when using Llama-3-8B-Instruct as the base model.
@@
Table 1: Results on AlpacaEval 2.0 and MT-Bench averaged with $3$ runs. Names inside the brackets are the models that are aligned based upon. The red arrows indicate the increment or decrement from the base model. Compared to iterative DPO and other online alignment baselines, SELM gains more improvements based on the weaker Zephyr-7B-SFT model and achieves superior performance that is competitive with much larger SOTA models when fine-tuned from Llama-3-8B-Instruct.
Notably, the implemented iterative DPO is obtained through comprehensive grid searches of hyperparameters and practical designs (see Appendix D for details), making it a strong baseline comparable with SOTA online alignment algorithms fine-tuned from more advanced models. For example, DPO Iter 3 (Zephyr) achieves an MT-Bench score of $7.46$, representing a $2.16$ improvement over Zephyr-SFT ($5.30$) and coming close to DNO ($7.48$), which is fine-tuned from the stronger model Orca-2.5-SFT ($6.88$). Additionally, SPPO achieves an MT-Bench score of $7.59$, a modest improvement of $0.08$ over Mistral-it ($7.51$). SELM leverages the optimal hyperparameters of iterative DPO while delivering improvements with almost zero additional computational overhead.
@@
Figure 2: Pairwise comparison between SELM, iterative DPO, and base models. Scores represent the LC win rates of the row models against the column models. Models positioned in higher rows have higher LC win rates against the base model and thus better performance.
Beyond instruction-following benchmarks, we also evaluate SELM and the baselines on several academic benchmarks, including GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), ARC challenge (Clark et al., 2018), TruthfulQA (Lin et al., 2021), EQ-Bench (Paech, 2023), and OpenBookQA (OBQA) (Mihaylov et al., 2018). To better reflect the capabilities of LLMs, we adopt various settings for these benchmarks, including zero-shot, few-shot, and few-shot Chain-of-Thought (CoT) settings. The accuracy results for these multiple-choice QA benchmarks are provided in Table 2. It can be observed that both our method and the baselines can degrade after the RLHF phase on some benchmarks, which is known as the alignment tax (Askell et al., 2021; Noukhovitch et al., 2024; Li et al., 2024b). Nevertheless, our method is still able to improve the base models on most of the benchmarks and offers the best overall performance.
We note that SELM is one of the instantiations of the proposed self-exploration objective in (1.1), with reparameterized reward functions and algorithm-specific designs described in Section 4.2, such as the dataset partition and update rule. However, this objective is not restricted to the current implementation and can also be directly applied to any other online alignment framework, with or without a separate reward model, regardless of differences in algorithm designs. Thus, the proposed method is orthogonal to and can be integrated directly into the recent online RLHF workflows (Dong et al., 2024; Xiong et al., 2023; Hu et al., 2024) that incorporate additional delicate designs with carefully curated datasets.
@@ 6.3 Ablation Study
Next, we study the difference in reward distributions with varied $\alpha$ and iterations. Specifically, for prompts from the 2k test set of UltraFeedback, we greedily sample from the LLM and generate rewards for the responses with PairRM. We then calculate the fraction of data that lies in each partition of rewards. The results for different $\alpha$ values of SELM Iter 2 (Zephyr) in Figure 3 (Middle) indicates that increasing $\alpha$ results in distributions that are concentrated in higher-reward regions.
@@
Figure 5: SELM actively explores by favoring high-reward responses.
In Section 5, we show that SELM engages in more active exploration by prioritizing high-reward responses compared to DPO, which indiscriminately favors unseen extrapolations and explores passively. To validate this, we sample three responses from SELM and DPO Iter 2 (Zephyr) for each prompt and we calculate the subtraction of the mean implicit rewards. As illustrated in Figure 5, SELM consistently achieves higher implicit rewards across most prompts, with the positive reward differences being notably larger in magnitude, supporting our claim regarding SELM’s active exploration behavior.
@@ 7 Conclusion & Future Work
In this paper, we introduced an active preference elicitation method for the online alignment of large language models. By incorporating an optimism term into the reward-fitting objective, the proposed bilevel self-exploring objective effectively balances between exploiting observed data and exploring potentially high-reward regions. Unlike standard online RLHF algorithms that passively explore the response space by sampling from the training LLM, whose sole objective is maximizing the expected learned reward, our method actively seeks diverse and high-quality responses. This self-exploration mechanism helps mitigate the risk of premature convergence and overfitting when the reward model is only locally accurate. To optimize this bilevel objective, we solve the inner-level problem and reparameterize the reward with the LLM policy, resulting in a simple yet novel iterative alignment algorithm called Self-Exploring Language Models (SELM). Compared to DPO, SELM is provably sample-efficient and improves the exploration efficiency by selectively favoring responses with high potential rewards rather than indiscriminately sampling unseen responses.
Our experiments, conducted with Zephyr-7B-SFT and Llama-3-8B-Instruct models, demonstrate the efficacy of SELM with consistent improvements on AlpacaEval 2.0, MT-Bench, and academic benchmarks with minimal computational overhead. These results underscore the ability of SELM to enhance the alignment and capabilities of LLMs by promoting more diverse and high-quality responses. Since the proposed technique is orthogonal to the adopted online RLHF workflow, it will be interesting to apply our method within more sophisticated alignment frameworks with advanced designs, which we would like to leave as future work.
@@
For the derivation of the DPO gradient $\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\mathcal{D}_{t})$, we refer the readers to Rafailov et al. (2024b).
@@
Appendix C Proof of Theorem 5.3
We use the reduction technique from Xie et al. (2024) to connect the sample complexity of SELM to that of existing RL algorithms (Zhong et al., 2022; Liu et al., 2024b). We restate the proof technique from Xie et al. (2024) for completeness. We emphasize that it is not a novel contribution of the present work. It is worth noting that the theoretical version of the self-exploration mechanism (Algorithm 2) is a bit different from the practical one used in the numerical experiments and is closer to the proposed algorithm in Xie et al. (2024).
We present the following theoretical version of the proposed self-exploration algorithm. The key modification in Algorithm 1 lies in its pragmatic strategy for constructing the chosen and rejected responses. Despite this adjustment, the core principles of leveraging the self-exploration objective during online alignment remain the same.
Algorithm 2 Self-Exploring Language Models (SELM; Theoretical Version)
0:  Reference model $\pi_{\text{ref}}$, preference dataset $\mathcal{D}_{0}=\varnothing$, prompt distribution $\nu$, online iterations $T$, optimism coefficient $\alpha$, $\pi_{0}=\pi_{\text{ref}}$.
1:  for iteration $t=1,2,\ldots,T$ do
2:     Sample $x_{t}\sim\nu$, $y_{t}^{1}\sim\pi_{t-1}(\cdot\mid x)$, $y_{t}^{2}\sim\pi_{\mathrm{ref}}(\cdot\mid x)$.
3:     Update the preference data $\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{(x_{t},y_{t}^{1},y_{t}^{2})\}$
4:     Train the LLM $\pi_{t}=\mathop{\mathrm{argmax}}_{\pi}\{-\mathcal{L}_{\text{DPO}}(\pi;\mathcal{D}_{t})-\alpha\cdot\mathbb{E}_{x\sim\nu}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}[\log\pi(y\mid x)]\}$, let $\pi_{\text{ref}}=\pi_{t}$.
5:  end for
Definition C.1 (Preference-based GEC).
For the function class $\Pi$, we define the preference-based GEC (PGEC) as the smallest $d_{\mathrm{GPEC}}$ as
$\sum_{t=1}^{T}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi_{t})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi_{t}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi_{t}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]$
$\leq\sqrt{d_{\mathrm{PGEC}}\sum_{t=1}^{T}\sum_{\tau=1}^{t-1}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi^{\tau})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{\tau}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi^{\tau}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]^{2}}$
$\qquad+4\sqrt{d_{\mathrm{PGEC}}T}.$
The definition of PGEC is a preference-based version of Generalized Eluder Coefficient (GEC) proposed by (Zhong et al., 2022). Intuitively, both PGEC and GEC establish a crucial connection between prediction error and in-sample estimation error, effectively transforming regret minimization into an online estimation problem. For a comprehensive explanation and in-depth discussion, readers are directed to Zhong et al. (2022). A slight difference is that the PGEC here is defined with respect to the policy class, while the GEC in Zhong et al. (2022) is defined with respect to the model or value class. These can be connected if we regard the implicit reward class $\log(\pi/\pi_{\mathrm{ref}})$ as the model or value class. As an important example, if we consider the log-linear function class $\Pi=\{\pi_{\theta}:\pi_{\theta}(y\mid x)\propto\exp(\langle\phi(x,y),\theta\rangle/\beta)\}$, we can show that $d_{\mathrm{PGEC}}=\widetilde{O}(d)$ by the elliptical potential lemma (Abbasi-Yadkori et al., 2011; Zhong et al., 2022). Another remark is that here the PGEC is defined in the bandit formulation, and it can be naturally extended to the token-wise MDP formulation (Zhong et al., 2024; Rafailov et al., 2024a; Xie et al., 2024) and further connects to the eluder dimension in the context of preference-based MDPs (Chen et al., 2022; Wang et al., 2023). Specifically, if we regard the generation process of LLMs as token-level MDPs where the generation of each token serves as one step, the learning objective is maximizing
$\mathcal{J}(\pi)=\mathbb{E}_{x\sim\nu,\tau\sim\pi}\left[r(\tau)-\beta\log\frac{\pi(\tau{\,|\,}x)}{\pi_{\mathrm{ref}}(\tau{\,|\,}x)}\right].$
Here $\tau$ is the full trajectory starting from $x$. We can similarly define the PGEC (Definition C.1) for token-wise MDPs by replacing the response $y,y^{\prime}$ in the bandit formulation with the trajectories $\tau,\tau^{\prime}$ in the token-wise MDP formulation. We have the following informal theorem:
Theorem C.2 (Regret for MDP Formulation (informal)).
With proper parameter choice, it holds with probability at least $1-\delta$ that
$\mathcal{R}(T)$ $\lesssim\sqrt{d_{\mathrm{PGEC}}\cdot\exp(2V_{\max})\cdot T\cdot\log(|\Pi|/\delta)},$
where $V_{\max}$ is a bounded coefficient for toekn-wise MDPs, similar to the one described in Assumption 5.2.
C.1 Proof of Theorem 5.3
Proof of Theorem 5.3.
We first decompose the regret as
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}[\mathcal{J}(\pi^{*})-\mathcal{J}(\pi_{t})]$
$=\sum_{t=1}^{T}\left(\mathbb{E}_{x\sim\nu,y\sim\pi^{*}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]-\mathbb{E}_{x\sim\nu,y\sim\pi_{t}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi_{t}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]\right)$
$=\sum_{t=1}^{T}\left(\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]-\mathbb{E}_{x\sim\nu,y\sim\pi_{t}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi_{t}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]\right),$
where the last line uses the fact that
$r(x,y)-\beta\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}=\beta\cdot\log Z_{r}(x),$ (C.1)
which is independent of the response $y$. Rearranging the above regret decomposition, we have
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\left(\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi^{*}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]-\mathbb{E}_{x\sim\nu,y\sim\pi_{t}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi_{t}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right]\right)$
$=\sum_{t=1}^{T}\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\left[\beta\log\frac{\pi_{t}(y\mid x)}{\pi^{*}(y\mid x)}\right]$
$\qquad+\sum_{t=1}^{T}\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x),y^{\prime}\sim\pi_{t}(\cdot\mid x)}\left[r(x,y)-\beta\log\frac{\pi_{t}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-r(x,y^{\prime})+\beta\log\frac{\pi_{t}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]$
$=\sum_{t=1}^{T}\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\left[\beta\log\frac{\pi_{t}(y\mid x)}{\pi^{*}(y\mid x)}\right]$
$\qquad+\beta\sum_{t=1}^{T}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi_{t})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi_{t}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi_{t}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right],$ (C.2)
where the last line uses (C.1). By the definition of PGEC in Definition C.1, we have
$\sum_{t=1}^{T}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi_{t})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi_{t}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi_{t}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]$
$\leq\sqrt{d_{\mathrm{PGEC}}\sum_{t=1}^{T}\sum_{\tau=1}^{t-1}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi^{\tau})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{\tau}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi^{\tau}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]^{2}}$
$\qquad+4\sqrt{d_{\mathrm{PGEC}}T}$
$\leq\frac{d_{\mathrm{PGEC}}}{4\eta}+\eta\sum_{t=1}^{T}\sum_{\tau=1}^{t-1}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi^{\tau})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{\tau}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi^{\tau}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]^{2},$
$\qquad+4\sqrt{d_{\mathrm{PGEC}}T},$ (C.3)
where the last inequality follows from the fact that $\sqrt{xy}\leq x/(4\eta)+\eta y$ for any $x,y,\eta>0$.
By the updating rule of $\pi_{t+1}=\mathop{\mathrm{argmax}}_{\pi}\{-\mathcal{L}_{\text{DPO}}(\pi;\mathcal{D}_{t})-\alpha\cdot\mathbb{E}_{x\sim\nu}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}[\log\pi(y\mid x)]\}$, we have
$-\mathcal{L}_{\text{DPO}}(\pi_{t};\mathcal{D}_{t-1})-\alpha\cdot\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}[\log\pi_{t}(y\mid x)]$
$\qquad\geq-\mathcal{L}_{\text{DPO}}(\pi^{*};\mathcal{D}_{t-1})-\alpha\cdot\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}[\log\pi^{*}(y\mid x)],$
which equivalents to that
$\mathbb{E}_{x\sim\nu,y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\left[\beta\log\frac{\pi_{t}(y\mid x)}{\pi^{*}(y\mid x)}\right]\leq\frac{\beta}{\alpha}\cdot\left(\mathcal{L}_{\mathrm{DPO}}(\pi^{*};\mathcal{D}_{t-1})-\mathcal{L}_{\mathrm{DPO}}(\pi_{t};\mathcal{D}_{t-1})\right).$ (C.4)
We upper bound the right handsise of (C.4) via the following lemma.
Lemma C.3 (Concentration).
For any $t\in[T]$ and $10<\delta<$ 1, it holds with probability $1-\delta$ that
$\mathcal{L}_{\mathrm{DPO}}(\pi^{*};\mathcal{D}_{t-1})-\mathcal{L}_{\mathrm{DPO}}(\pi_{t};\mathcal{D}_{t-1})$
$\lesssim-\frac{2}{\exp(4R_{\max})}\cdot\sum_{\tau=1}^{t-1}\mathbb{E}_{(x,y,y^{\prime})\sim(\nu,\pi_{\mathrm{ref}},\pi^{\tau})}\left[\log\frac{\pi^{*}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{\tau}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}-\log\frac{\pi^{*}(y^{\prime}{\,|\,}x)}{\pi_{\mathrm{ref}}(y^{\prime}{\,|\,}x)}+\log\frac{\pi^{\tau}(y^{\prime}\mid x)}{\pi_{\mathrm{ref}}(y^{\prime}\mid x)}\right]^{2}$
$\qquad+\log(|\Pi|/\delta).$
Proof.
The proof of this lemma follows the standard MLE analysis (Zhang, 2006) and its application for standard reward-based RL (Agarwal et al., 2020; Liu et al., 2024b). Recent works (Liu et al., 2024c; Xie et al., 2024; Cen et al., 2024) also applies this result for RLHF. For brevity, we omit the detailed proof here and direct readers to these related works for the proof. ∎
Combining (C.1), (C.1), (C.4), and Lemma C.3, together with the parameter choice $\alpha=2/(\eta\exp(4R_{\max}))$, we obtain
$\mathcal{R}(T)$ $\lesssim\frac{\beta Td_{\mathrm{PGEC}}}{\eta}+\beta\eta\cdot{\exp(4R_{\max})}\log(|\Pi|/\delta)+4\sqrt{d_{\mathrm{PGEC}}T}$
$\lesssim\sqrt{d_{\mathrm{PGEC}}\cdot\exp(2R_{\max})\cdot T\cdot\log(|\Pi|/\delta)},$
where the last line follows from the fact that $\eta=\sqrt{Td_{\mathrm{PGEC}}/(\exp(4R_{\max})\log(|\Pi|/\delta))}$. Therefore, we finish the proof of Theorem 5.3. ∎
Appendix D Experiment Setup